A country has 13,000 hours of labor to allocate across the production of two types of products: iPhones and apples. Each iPhone requires 30 hours of labor to produce and each apple requires 40 hours of labor to produce. In a world with scarce food and hungry technology consumers, surprisingly both goods sell for $100 per unit on the open market.
The country's policy makers are considering two options for production:
Option A: 200 iPhones and 150 apples
Option B: 100 iPhones and 250 apples
Which of the following is true?

Both options are feasible, but only Option A is efficient.
Options A and B are equally preferred since they both produce the same total number of products.
Option B is preferred since it is feasible and efficient while Option A is not efficient.
Options A and B are equally preferred since they're both feasible.
Option A is preferred since it is feasible and efficient while Option B is not efficient.
Answer #1

The country has a total of 13,000 labor hours. Since both goods are sold at the same price, we do not create a distinction based on the revenue generated because a total of 350 {(200,150), (100, 250)} units are being sold at same price across both the goods. That is, the revenue remains constant at any option.

Total Revenue for both options : $100 x 350 = $350,000

We check for the feasibility (the total labor hours used being less than or equal to 13,000) and efficiency (not wasting labour hours that is available, utilizing the labor resource fully) of the two options.

Option A : 200 iPhones => 30 x 200 = 6,000 labor hours ; 150 apples => 40 x 150 = 6,000 labor hours

Total labor hours utlized = 6,000 + 6,000 = 12,000

Option B : 100 iPhones => 30 x 100 = 3,000 labor hours; 250 apples => 40 x 250 = 10,000 labor hours

Total labor hours utlized = 3,000 + 10,000 = 13,000

Both options are feasible because each is utilizing under 13,000 labor hours. But option A is not efficient. It is leaving 1,000 labor hours left unused and not fully utilizing labor resource available while generating similar total revenue as that in option B.

Correct Answer : Option B is preferred since it is feasible and efficient while Option A is not efficient.